# Training Hybrid Deep Quantum Neural Networks for Efficient Reinforcement Learning

### v5

@@ A.1.2 The inefficiency of training hDQNN without qtDNN surrogate
Because $I\times O$ different circuits are required to estimate all the gradients of a QL with $I$ input dimensions and $O$ output dimensions and each circuit needs a number, $S$, of shots to have the statistics for estimating the observables used to calculate gradients. Moreover, learning for complex problems with a high sample data efficiency would require training data to be sampled from past experiences in mini-batches of size $N_{\text{b}}$ for each DNN update with batch optimization as in Haarnoja et al. [2018]. Therefore, existing QML methods that directly estimate gradients of the QL are faced with a prohibitive practical training overhead proportional to $I\times O\times S\times N_{\text{b}}$ for running on real QPU that can only run one shot at a time. For example, in the case of using a fast superconducting Josephson junction quantum processor that has gate time as short as 20 ns and measurement time in the range of 100 ns as seen in Acharya et al. [2025], Gao et al. [2025], a 10-layer 10-qubit PQC (QL) would need $I=220$ and $O=10$ with $N_{\text{b}}=256$ and $S=1000$ in our example for learning complex Humanoid-v4 environment. In this case, each update to model DNN parameters would require about 281.6 s. Considering that typical classical RL training even for simple tasks would likely require $\sim 10^{6}$ update steps similar to Haarnoja et al. [2018], this would translate to more than 10 years of training.
A.2 Proof of Theorem  1 on the Local gradient fidelity of qtDNN
@@
Because $J_{Q}$ is $L$-Lipschitz and $\|h\|\leq r$, from (A.2) we obtain
[…]
Differentiating (A.2) with respect to $x$ yields, using $\nabla_{x}h=I$,
[…]
@@
For any $x\in\mathcal{B}$, combine (A.2) with (A.2):
@@
Likewise, using (A.2) and (A.2),
@@
Appendix B Mathematical formalization of the hDQNN-TD3’s components
The parameterized quantum circuit (PQC) in the hDQNN-TD3 is a variational circuit designed for NISQ devices, utilizing $N$ qubits to act as a quantum hidden layer within the actor network, taking its inputs from the DNN’s previous layer’s extracted features. It consists of $M$ layers, each comprising single-qubit rotation gates and entangling CZ gates.
Formally, the PQC transforms an input vector $\mathbf{q_{i}}$ into a probability distribution over bitstrings $\mathbf{q_{o}}$ by the following:
@@ Appendix B Mathematical formalization of the hDQNN-TD3’s components
When measuring the qubits in the computational basis, one obtains a bitstring $\mathbf{q_{o}}\in\{0,1\}^{N}$, with probability:
@@
B.1 Equations for the qtDNN surrogate
@@
During each forward pass the real PQC (or its CUDA–Q simulation) is executed to generate a bit-string $\mathbf{q_{o}}\in\{0,1\}^{N}$; the qtDNN is invoked only in the backward pass as a differentiable proxy that supplies $\nabla_{\theta}\hat{\mathbf{p}}$ without additional quantum evaluations. To this end we draw minibatches of pairs $\bigl{(}\mathbf{q_{i}},\mathbf{q_{o}}\bigr{)}$ where $\mathbf{q_{i}}\in\mathbb{R}^{2NM}\bigotimes\mathbb{N}^{2M}$ are the rotation angles and pairs of qubit indices fed to the PQC and $\mathbf{q_{o}}\in\{0,1\}^{N}$ is a bit-string obtained from measurements over multiple shots of executing the same quantum circuit. In hDQNN-TD3, the bit-string is taken as the most probable string among the bit-strings obtained from multiple shots. The qtDNN learns a mapping
@@
Ideally one would minimise the cross-entropy with respect to the joint distribution produced by the PQC, $P_{\text{PQC}}(\mathbf{q_{o}}\mid\boldsymbol{\theta})$, namely $\mathcal{L}=-\log P_{\text{PQC}}(\mathbf{q_{o}}\mid\boldsymbol{\theta})$. However estimating that joint requires $\Omega(2^{N})$ shots per input and is infeasible even for the moderate $N=10$ qubits used in our study. Following prior work on quantum-model surrogates, we therefore adopt the factorised binary cross-entropy (BCE)
@@
$=\;\mathbb{E}_{(\boldsymbol{\theta},\mathbf{q}_{\text{o}})\sim\mathcal{M}}\bigl{[}\operatorname{BCE}\!\bigl{(}\mathbf{q_{o}},f_{\text{qtDNN}}(\mathbf{q_{i}};\omega)\bigr{)}\bigr{]},$ (6)
where $\mathcal{M}$ is the minibatch of replay buffer of observed $\bigl{(}\mathbf{q_{i}},\mathbf{q_{o}}\bigr{)}$ pairs.
@@
Equation (5) treats qubits as independent Bernoulli variables, whereas the true PQC may entangle them. We demonstrated theoretically that this approximation is adequate for gradient propagation. Moreover, the stabtility of the learning process shows that the surrogate successfully reproduces the expectation values of the PQC. Future work could replace the factorised model with an autoregressive density estimator to capture higher-order correlations without the exponential shot overhead.
B.2 Forward pass of the hDQNN (inference, no RL)
@@
Given an observation $\mathbf{x}\in\mathbb{R}^{n}$, the PreDNN (a feed-forward MLP) produces a latent vector $\mathbf{z}_{\mathrm{pre}}\in\mathbb{R}^{(2N+2)M+d_{\text{c-link}}}$, where the last $d_{\text{c-link}}$ elements are passed to PostDNN directly. The rest of the elements are rescaled to $[{-}\pi,\pi]^{\otimes 2NM}$ or rounded to integers $\mathbb{N}^{2M}$ as follows. The rescaled part is split into the pair of rotation angles for every qubit in every layer of the PQC,
$\begin{split}\boldsymbol{\theta}&=\mathbf{z}_{\mathrm{pre}}^{1{:}N\times M},\qquad\boldsymbol{\phi}=\mathbf{z}_{\mathrm{pre}}^{N\times M+1{:}2N\times M},\qquad\boldsymbol{k}=\mathbf{z}_{\mathrm{pre}}^{2N\times M+1{:}(2N+2)M},\\ \mathbf{q}_{\mathrm{i}}&=(\boldsymbol{\theta},\boldsymbol{\phi},\boldsymbol{k})\in[-\pi,\pi]^{2N\times M}\otimes\mathbb{N}^{2M}.\end{split}$ (7)
@@
B.3 Backward pass with the qtDNN surrogate (supervised training, no RL)
@@
The loss $\mathcal{L}_{\text{sup}}$ is task-specific (e.g. MSE or cross-entropy). Because $f_{\text{qtDNN}}$ is an ordinary neural network, automatic differentiation supplies $\nabla_{\theta_{\text{Post}}}\mathcal{L}_{\text{sup}}$, $\nabla_{\omega}\mathcal{L}_{\text{sup}}$, and $\nabla_{\theta_{\text{Pre}}}\mathcal{L}_{\text{sup}}$ without invoking the quantum simulator. During the same update step we refine the surrogate itself by a second objective, the BCE in Eq. (5), computed on the cached pair $(\mathbf{q}_{\mathrm{i}},\mathbf{z}_{\mathrm{q}})$ from the forward PQC evaluation:
@@
B.4 Reinforcement Learning
The hDQNN-TD3’s forward pass with RL is already included in the main paper, in Algorithm 1. This section will simply introduce the underlying concepts in higher depth.
B.4.1 Notations and Definitions
@@
[…]
@@
B.4.2 Actor-Critic Framework
@@
    Let $C:(s,\,a)\mapsto\hat{R}$ be the critic, parameterized by $\theta_{C}$. It estimates the expected cumulative reward (or return). We define the one-step temporal-difference (TD) target:
@@
    $\mathbf{q_{o}}$ $\sim P\bigl{(}\mathbf{q_{o}}\mid\mathbf{q_{i}}\bigr{)}\;=\;\text{(Measurement of PQC parameterized by }\mathbf{q_{i}}),$
@@
    where $\mathbf{z}_{\text{c-link}}$ indicates possible direct links from the output of $f_{\text{preDNN}}$ to the input of $f_{\text{postDNN}}$. The distribution $P(\mathbf{q_{o}}\mid\mathbf{q_{i}})$ captures the probability of measuring a particular bitstring $\mathbf{q_{o}}$ on the quantum hardware.
@@
    In practice, gradients $\nabla_{\theta_{A}}\mathcal{L}_{A}$ are computed. Here, the qtDNN can be used to approximate the map $\mathbf{q_{i}}\mapsto\mathbf{q_{o}}$ locally. This allows batched backpropagation through the classical layers without repeatedly querying the quantum circuit for gradients.
B.5 Forward pass of hDQNN–TD3 (interaction phase)
@@
[…]
@@
B.6 Backward pass of hDQNN–TD3 (update phase)
Every $u$ environment steps we draw a minibatch $\mathcal{B}\subset\mathcal{D}$ and perform the following updates.

### v6

@@ A.1.2 The inefficiency of training hDQNN without qtDNN surrogate
Because $I\times O$ different circuits are required to estimate all the gradients of a QL with $I$ input dimensions and $O$ output dimensions and each circuit needs a number, $S$, of shots to have the statistics for estimating the observables used to calculate gradients. Moreover, learning for complex problems with a high sample data efficiency would require training data to be sampled from past experiences in mini-batches of size $N_{\text{b}}$ for each DNN update with batch optimisation as in Haarnoja et al. [2018]. Therefore, existing QML methods that directly estimate gradients of the QL are faced with a prohibitive practical training overhead proportional to $I\times O\times S\times N_{\text{b}}$ for running on real QPU that can only run one shot at a time. For example, in the case of using a fast superconducting Josephson junction quantum processor that has gate time as short as 20 ns and measurement time in the range of 100 ns as seen in Acharya et al. [2025], Gao et al. [2025], a 10-layer 10-qubit PQC (QL) would need $I=220$ and $O=10$ with $N_{\text{b}}=256$ and $S=1000$ in our example for learning complex Humanoid-v4 environment. In this case, each update to model DNN parameters would require about 281.6 s. Considering that typical classical RL training even for simple tasks would likely require $\sim 10^{6}$ update steps similar to Haarnoja et al. [2018], this would translate to more than 10 years of training.
A lighter but still impractical parameter‑shift budget.
Consider an optimistic configuration with $S=10$ shots per circuit, $N_{b}=256$, $I=220$ inputs and $O=10$ outputs for the N=10,M=10N{=}10,M{=}10 PQC, and single‑/two‑qubit/measurement latencies of $(50\,\mathrm{ns},50\,\mathrm{ns},100\,\mathrm{ns})$. Each actor update requires $I\times O\times S\times N_{b}\approx 5.63\times 10^{6}$ shots, i.e. $\approx 3.38\,\mathrm{s}$ at $\sim 0.6\,\mu\mathrm{s}$ per shot. Over $\sim 7.5\times 10^{5}$ updates (7 200 episodes), this exceeds $2.5\times 10^{6}\,\mathrm{s}$ ($\approx 29$ days) before queueing/calibration overhead. By contrast, the qtDNN run in Table 2 completes in $\sim 700$ minutes.
A.1.3 qtDNN versus classical simulators.
A noiseless state-vector simulator stores $2^{N}$ amplitudes and updates them gate-by-gate, with time/memory complexity $O(2^{N}L)$. Even modest circuits thus exceed the footprint required for parallel mini-batches. More importantly, gradients from an ideal simulator would ignore read-out error, decoherence, and cross-talk present at inference on NISQ QPUs. In contrast, qtDNN is fitted on the noisy measurements $\mathbf{q}_{\text{o}}$ at every step, so the gradients used to train the actor are aligned with the behaviour that the policy will experience at deployment.
A.1.4 Complexity of qtDNN updates
One surrogate fit touches $W$ weights, hence costs $O(W)$ FLOPs per tiny-batch. In our 10-qubit setup with PostDNN bypass of $d_{\text{c-link}}=10$ and a single hidden layer $L_{h}=2048$, the surrogate has $W\approx 2048\,(2N{+}2)M+2048\,10+2048\cdot 10\approx 4.7\times 10^{5}$ parameters ($\approx 1.8$ MB in FP32). By contrast, an exact state-vector simulator stores $2^{N}$ complex amplitudes and updates them per gate, $O(2^{N}L)$, while a parameter-shift estimator incurs $\Theta(I)$ additional circuit evaluations per input dimension, returning to $O(I\,2^{N}L)$ time and memory. This explains the 2–3 orders of magnitude gap between our qtDNN updates and simulator-based gradients at the N=10N{=}10, L=10L{=}10 scale reported in Table 3.
Appendix B Local gradient fidelity
B.1 Proof of Theorem  1 on the Local gradient fidelity of qtDNN
Summary of proof.
Because the mini-batch $\mathcal{B}$ lies inside a small radius-$r$ ball, $Q$ is well-approximated by its first-order Taylor expansion around the centre $\bar{x}$. Thus, we (i) approximate the first‑order Taylor polynomial $P(x){:=}Q(\bar{x}){+}J_{Q}(\bar{x})(x{-}\bar{x})$ over $\mathcal{B}$ by an MLP $Q_{\theta}$, and (ii) use Lipschitz bounds on the second derivative of $Q$ to control the Taylor remainder and its gradient. Replacing the PQC by that MLP therefore changes (i) forward activations by at most $\varepsilon_{1}$ and (ii) their Jacobian by at most $\varepsilon_{2}$, which in turn perturbs the gradient flowing into earlier layers by the same order of magnitude. See below for the full proof.
@@ B.1 Proof of Theorem  1 on the Local gradient fidelity of qtDNN
Because $J_{Q}$ is $L$-Lipschitz and $\|h\|\leq r$, from equation B.1 we obtain
$\|\rho(x)\|\leq\int_{0}^{1}\!L\,t\,\|h\|^{2}\,dt=\frac{1}{2}L\|h\|^{2}\leq\frac{1}{2}Lr^{2}.$
Differentiating equation B.1 with respect to $x$ yields, using $\nabla_{x}h=I$,
$\nabla\rho(x)=\int_{0}^{1}t\nabla^{2}Q(\bar{x}+th)\,hdt.$
@@
For any $x\in\mathcal{B}$, combine equation B.1 with equation B.1:
@@
Likewise, using equation B.1 and equation B.1,
@@
B.2 Smoothness of PQC expectations.
In our hardware‑efficient ansatz all tunable gates are smooth rotations $R_{\alpha}(\theta)=\exp(-i\theta\,\sigma_{\alpha}/2)$ and entanglers are fixed; composition and the exponential map are analytic. Therefore the unitary $U(\bm{\theta})$ and the expectations $\langle Z_{k}\rangle(\bm{\theta})=\langle 0|U^{\dagger}(\bm{\theta})Z_{k}U(\bm{\theta})|0\rangle$ are $C^{\infty}$ in $\bm{\theta}$, which justifies the hypothesis $Q\in C^{1}$ used in Theorem 1. If one inserts non‑smooth operations (e.g., hard thresholds or topology‑switching branches), $Q$ may only be piecewise differentiable; in that regime the tight mini‑batch gradient fidelity does not apply verbatim and one must resort to sub‑gradient or sampling‑based estimators with typically higher variance.
B.3 Refitting vs freezing the surrogate.
After each actor update step the policy parameters change, shifting the distribution of PQC inputs seen at the next update. Holding the surrogate fixed across epochs would increase its BCE on the new batch, risking violation of the $\varepsilon$-ball required by Theorem 1. Because a qtDNN fit accounts for only $\sim 2\%$ of wall‑clock on our A100 runs (App. G.3), we refit every actor update and freeze the surrogate only for the remainder of that update epoch.
Appendix C Mathematical formalisation of the hDQNN-TD3’s components
The parameterised quantum circuit (PQC) in the hDQNN-TD3 is a variational circuit designed for NISQ devices, utilizing $N$ qubits to act as a quantum hidden layer within the actor network, taking its inputs from the DNN’s previous layer’s extracted features. It consists of $M$ layers, each comprising single-qubit rotation gates and entangling CZ gates.
Formally, the PQC transforms an input vector $\mathbf{q_{i}}$ into a probability distribution over bit-strings $\mathbf{q_{o}}$ by the following:
@@ Appendix C Mathematical formalisation of the hDQNN-TD3’s components
When measuring the qubits in the computational basis, one obtains a bit-string $\mathbf{q_{o}}\in\{0,1\}^{N}$, with probability:
@@
C.1 Equations for the qtDNN surrogate
@@
During each forward pass the real PQC (or its CUDA–Q simulation) is executed to generate a bit-string $\mathbf{q_{o}}\in\{0,1\}^{N}$; the qtDNN is invoked only in the backward pass as a differentiable proxy that supplies $\nabla_{\theta}\hat{\mathbf{p}}$ without additional quantum evaluations. To this end we draw mini-batches of pairs $\bigl(\mathbf{q_{i}},\mathbf{q_{o}}\bigr)$ where $\mathbf{q_{i}}\in\mathbb{R}^{2NM}\bigotimes\mathbb{N}^{2M}$ are the rotation angles and pairs of qubit indices fed to the PQC and $\mathbf{q_{o}}\in\{0,1\}^{N}$ is a bit-string obtained from measurements over multiple shots of executing the same quantum circuit. In hDQNN-TD3, the bit-string is taken as the most probable string among the bit-strings obtained from multiple shots. The qtDNN learns a mapping
@@
Ideally one would minimise the cross-entropy with respect to the joint distribution produced by the PQC, $P_{\text{PQC}}(\mathbf{q_{o}}\mid\bm{\theta})$, namely $\mathcal{L}=-\log P_{\text{PQC}}(\mathbf{q_{o}}\mid\bm{\theta})$. However estimating that joint requires $\Omega(2^{N})$ shots per input and is infeasible even for the moderate $N=10$ qubits used in our study. Following prior work on quantum-model surrogates, we therefore adopt the factorised binary cross-entropy (BCE)
@@
$=\;\mathbb{E}_{(\bm{\theta},\mathbf{q}_{\text{o}})\sim\mathcal{M}}\bigl[\operatorname{BCE}\!\bigl(\mathbf{q_{o}},f_{\text{qtDNN}}(\mathbf{q_{i}};\omega)\bigr)\bigr],$ (6)
where $\mathcal{M}$ is the mini-batch of replay buffer of observed $\bigl(\mathbf{q_{i}},\mathbf{q_{o}}\bigr)$ pairs.
@@
Equation equation 5 treats qubits as independent Bernoulli variables, whereas the true PQC may entangle them. We demonstrated theoretically that this approximation is adequate for gradient propagation. Moreover, the stability of the learning process shows that the surrogate successfully reproduces the expectation values of the PQC. Future work could replace the factorised model with an autoregressive density estimator to capture higher-order correlations without the exponential shot overhead.
C.2 Forward pass of the hDQNN (inference, no RL)
@@
Given an observation $\mathbf{x}\in\mathbb{R}^{n}$, the PreDNN (a feed-forward MLP) produces a latent vector $\mathbf{z}_{\mathrm{pre}}\in\mathbb{R}^{(2N+2)M+d_{\text{c-link}}}$, where the last $d_{\text{c-link}}$ elements are passed to PostDNN directly. The rest of the elements are rescaled to $\pi,\pi]^{\otimes 2NM}$ or rounded to integers $\mathbb{N}^{2M}$ as follows. The rescaled part is split into the pair of rotation angles for every qubit in every layer of the PQC,
$\begin{split}\bm{\theta}&=\mathbf{z}_{\mathrm{pre}}^{1{:}N\times M},\qquad\bm{\phi}=\mathbf{z}_{\mathrm{pre}}^{N\times M+1{:}2N\times M},\qquad\bm{k}=\mathbf{z}_{\mathrm{pre}}^{2N\times M+1{:}(2N+2)M},\\ \mathbf{q}_{\mathrm{i}}&=(\bm{\theta},\bm{\phi},\bm{k})\in[-\pi,\pi]^{2N\times M}\otimes\mathbb{N}^{2M}.\end{split}$ (7)
@@
C.3 Backward pass with the qtDNN surrogate (supervised training, no RL)
@@
The loss $\mathcal{L}_{\text{sup}}$ is task-specific (e.g. MSE or cross-entropy). Because $f_{\text{qtDNN}}$ is an ordinary neural network, automatic differentiation supplies $\nabla_{\theta_{\text{Post}}}\mathcal{L}_{\text{sup}}$, $\nabla_{\omega}\mathcal{L}_{\text{sup}}$, and $\nabla_{\theta_{\text{Pre}}}\mathcal{L}_{\text{sup}}$ without invoking the quantum simulator. During the same update step we refine the surrogate itself by a second objective, the BCE in Eq. equation 5, computed on the cached pair $(\mathbf{q}_{\mathrm{i}},\mathbf{z}_{\mathrm{q}})$ from the forward PQC evaluation:
@@
C.4 Reinforcement Learning
The hDQNN-TD3’s forward pass with RL is already included in App. D, as Algorithm 2. This section will simply introduce the underlying concepts in higher depth.
C.4.1 Notations and Definitions
@@
    $s$ $\in\mathbb{R}^{n_{s}}\quad\text{(the observed state in practice)},$
    $a$ $\in\mathbb{R}^{n_{a}}\quad\text{(the action suggested by the actor)}.$
@@
C.4.2 Actor-Critic Framework
@@
    Let $C:(s,\,a)\mapsto\hat{R}$ be the critic, parameterised by $\theta_{C}$. It estimates the expected cumulative reward (or return). We define the one-step temporal-difference (TD) target:
@@
    $\mathbf{q_{o}}$ $\sim P\bigl(\mathbf{q_{o}}\mid\mathbf{q_{i}}\bigr)\;=\;\text{(Measurement of PQC parameterised by }\mathbf{q_{i}}),$
@@
    where $\mathbf{z}_{\text{c-link}}$ indicates possible direct links from the output of $f_{\text{preDNN}}$ to the input of $f_{\text{postDNN}}$. The distribution $P(\mathbf{q_{o}}\mid\mathbf{q_{i}})$ captures the probability of measuring a particular bit-string $\mathbf{q_{o}}$ on the quantum hardware.
@@
    In practice, gradients $\nabla_{\theta_{A}}\mathcal{L}_{A}$ are computed. Here, the qtDNN can be used to approximate the map $\mathbf{q_{i}}\mapsto\mathbf{q_{o}}$ locally. This allows batched back-propagation through the classical layers without repeatedly querying the quantum circuit for gradients.
C.5 Forward pass of hDQNN-TD3 (interaction phase)
@@
    Actor evaluation. The current observation $s_{t}$ is propagated through the hybrid actor exactly as in Sec. C.2, i.e. $s_{t}\!\xrightarrow{\text{PreDNN}}\!\mathbf{q}_{\mathrm{i},t}\!\xrightarrow{\text{PQC}}\!\mathbf{z}_{\mathrm{q},t}\!\xrightarrow{\text{PostDNN}}\!a_{t}$. Exploration noise $\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ is then added:
@@
C.6 Backward pass of hDQNN-TD3 (update phase)
Every $u$ environment steps we draw a mini-batch $\mathcal{B}\subset\mathcal{D}$ and perform the following updates.